Fairbanks Corporation purchased 400 shares of Sherman Inc

Brief Exercise 17-05 


Fairbanks Corporation purchased 400 shares of Sherman Inc. common stock for $13,200 (Fairbanks does not have significant influence). During the year, Sherman paid a cash dividend of $3.25 per share. At year-end, Sherman stock was selling for $34.50 per share. 


Prepare Fairbanks's journal entries to record (a) the purchase of the investment, (b) the dividends received, and (c) the fair value adjustment. (Assume a zero balance in the Fair Value Adjustment account.) (Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter o for the amounts.) 


No. Account Titles and Explanation    Debit    Credit 

(a)

(b)

(c)
Answer #1

Ans:

SL NO ACCOUNT TITLE DEBIT CREDIT
A EQUITY INVESTMENT $13200
CASH $13200
B CASH $1300
DIVIDEND REVENUE $1300
C EQUITY INVESTMENT $600
UNREALIZED HOLDING GAIN OR LOSS-EQUITY $600

Notes:

A.$13200-given

B.$3.25*400=$1300

C.Fair value -Purchase price=(400*34.50)-13200=$600